OpenStudy (anonymous):

what changes we make in different quadrants to find argument..give the difference 1)Convert the given complex number in polar form: 1 – i 2)Convert the given complex number in polar form: – 1 + i

OpenStudy (anonymous):

@phi

OpenStudy (phi):

do you know trigonometry?

OpenStudy (anonymous):

yes ofcourse

OpenStudy (phi):

plot 1 – i and find the angle (measured from the x-axis, going counter-clockwise)

OpenStudy (anonymous):



OpenStudy (phi):

next, find the length of the hypotenuse in this case sqr(2) now write \[ Ae^{i\theta} \] where A is the length and theta is the angle

OpenStudy (phi):

people generally write the angle in radians

OpenStudy (phi):

\[ 1 – i = \sqrt{2}e^{-i \frac{\pi}{4}} \] \[ -1 + i = \sqrt{2}e^{i \frac{3\pi}{4}} \] notice if we add pi to the first number's angle, we get 3pi/4 and if we look at the plot we see the two points form a line.

OpenStudy (anonymous):

\[\cos \theta=1/\sqrt{2} and \sin \theta=-1/\sqrt{2}\] i have got this so far now i want to know it further

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@goformit100

OpenStudy (phi):

I posted the result up above \[ 1 – i = \sqrt{2}e^{-i \frac{\pi}{4}} = \sqrt{2}\left(\cos\left( -\frac{\pi}{4}\right) + i \sin\left( -\frac{\pi}{4}\right)\right) \] using trig identities we can write that as \[ \sqrt{2}\left(\cos\left( \frac{\pi}{4}\right) - i \sin\left( \frac{\pi}{4}\right)\right) \] note that if we sub in the numeric values for cos and sin we get \[ \sqrt{2}\left(\frac{1}{\sqrt{2}} - i \frac{1}{\sqrt{2}}\right) = 1 - i\] and we recover the rectangular form

OpenStudy (anonymous):

cant it be like sintheta/costheta=tantheta

OpenStudy (anonymous):

coz we are taught like this and it is acceptable only

OpenStudy (anonymous):

\large \theta = \tan^{-1}(-1) = -\dfrac{\pi}{4}

OpenStudy (phi):

sin(x)/cos(x) = tan(x) is an identity Do you have a specific question? Perhaps you are using a different notation for the polar form? These all mean the same thing \[ r \ cis \ \theta \\ r \angle \theta \\ r e^{i \theta} \\ r(\cos \theta + i \sin \theta)\]

OpenStudy (phi):

to find theta associated with complex number x you would use \[ \tan^{-1}\left( \frac{imag(x)}{real(x)}\right) \]

OpenStudy (anonymous):

like we got r=\[\sqrt{2}\]

OpenStudy (phi):

I would use pythagorean theorem to find r \[ r = \sqrt{real(x))^2 + imag(x)^2} \]

OpenStudy (anonymous):

\[i use r(costheta+isintheta)\]

OpenStudy (phi):

then the answer can be written as \[\sqrt{2}\left(\cos\left( -\frac{\pi}{4}\right) + i \sin\left( -\frac{\pi}{4}\right)\right) \\ \sqrt{2}\left(\cos\left( \frac{\pi}{4}\right) - i \sin\left( \frac{\pi}{4}\right)\right)\\ \sqrt{2}\cos\left( \frac{\pi}{4}\right) - i \sqrt{2}\sin\left( \frac{\pi}{4}\right) \]

OpenStudy (anonymous):

but how do we get pi/4 ????????

OpenStudy (phi):

or \[ \sqrt{2} \left( \cos(\frac{7 \pi}{4}) + i \sin(\frac{7 \pi}{4}) \right)\]

OpenStudy (phi):